Suppose 5% of cereal boxes contain a prize. You are determined to buy cereal boxes until you win a prize.

A) What kind of probability distribution is this, and explain your choice.

B) What is the probability that you win a prize on the 2nd box?

C) Find the mean and standard deviation of this distribution
Answer #1

X: number of cereal boxes purchased until you win a prize.

a). we will use GEOMETRIC distribution.

[ here the number of trials are unknown .Also,  the trials are independent and the probability of success (p) is fixed.so, we can conclude that X ~ geometric (0.05 ).

the pmf of the distribution is :-

P(X=x)=(1-0.05)^{x-1}*0.05]

b). the probability that you win a prize on the 2nd box is:-

P(X=2)=(1-0.05)^{2-1}*0.05=\mathbf{0.0475}

c).the mean of this distribution is :-

=\frac{1}{p}=\frac{1}{0.05}=\mathbf{20}

standard deviation of the distribution is :-

=\sqrt{\frac{1-p}{p^2}}=\sqrt{\frac{1-0.05}{0.05^2}}\approx\mathbf{19.49}

*** if you have any doubt regarding the problem please write it in the comment box.if you are satisfied please give me a LIKE if possible...